### v5

# Impact of chaotic magnetic field on mass-radius relation of rotating neutron stars

Abstract
Observations reveal that magnetic fields on neutron stars (NSs) are in the range of $10^{8-15}$ G. Apart from being celestial bodies, NSs are normally rotating. In this work, we study the impact of a chaotic magnetic field on the mass-radius relation of the rotating NSs. We employ an equation of state of NSs with the nuclei in the crust and hyperons in the core. We use Hartle-Thorne formalism as an approximation of the rotating NSs. For the magnetic field ansatz, we use the one coupled to the energy density. We find that the magnetic field can decrease radius of NS. NSs formed with stronger chaotic magnetic fields exhibit a lower maximum mass compared to those formed with weaker chaotic magnetic fields. In contrast, the increment of the magnetic field can increase the compactness and deformation of rotating NSs.
1 Introduction
It has been widely understood that the evolution of the massive stars whose mass is 8--25 $M_{\odot}$ ended when neutron stars (NSs) formed. The corresponding formation is due to the stars’ gravitational collapse during Type-II, Ib, or Ic supernova explosion phenomena [1, 2]. NSs are dense, compact remnants of evolved stars, with degenerate fermions compressed by strong gravitational fields [3]. NSs are normally rotating as all celestial objects tend to naturally always rotate due to the law of angular momentum conservation [4]. Rapidly rotating NSs are observed as pulsars [5].
NSs have a wide range of magnetic fields. Observations indicate that magnetic fields on NSs are at least in the range of $10^{8-15}$ G. New classes of pulsars such as Anomalous X-ray pulsars (AXPs) and Soft-Gamma Repeater pulsars (SGRs) have been identified producing vast magnetic fields. SGR is associated with remnants of a supernova, which is a young NS [6]. Highly magnetized NSs, i.e. magnetars, are typically slowly spinning NSs whose emission is thought to be powered by the decay of their large ($\geq 10^{12}$ G) magnetic fields [7]. Moreover, observations on some AXPs also indicate that AXPs surface magnetic field is around $10^{14}-10^{15}$ G [6].
Konno et al. [8] calculated the deformation of NSs resulting from rotation and the presence of magnetic fields. In their analysis, a polytropic equation of state (EOS) was employed. Their findings indicate that for magnetars, the deformation induced by magnetic fields is dominant, whereas for typical pulsars, the magnetic effect is negligible. Mallick & Schramm [9] investigated mass corrections and the deformation of NSs under the influence of magnetic fields, treating the magnetic field as an anisotropy in the energy-momentum tensor (EMT). Interestingly, despite focusing on static NSs, their approach was based on the Hartle-Thorne (HT) formalism [10, 11], which is typically applied to rotating relativistic stars. In their work, magnetic perturbations in static configurations were treated analogously to rotational perturbations, following the HT approximation. They considered both very stiff and very soft EOS models in their study. Lopes & Menezes [12] explored the effects of chaotic magnetic fields on NSs. A key advantage of incorporating a chaotic magnetic field lies in the elimination of anisotropy, simplifying the mathematical formulation. They also introduced an ansatz where the magnetic field is coupled to the energy density of the NS. For their EOS models, they utilized two variations: one including hyperons and one without. However, the rotation of NSs was not considered in their analysis.
Building on previous studies, we investigate the influence of the chaotic magnetic field on the mass-radius relationship of rotating NSs. The rotational aspects are modeled using the HT formalism, while the magnetic field is described by the Lopes-Menezes ansatz. The EOS employed in this work is based on the realistic model proposed by Miyatsu et al. [13], which includes nuclei in the crust and hyperons in the core. The algorithm of the numerical simulation follows the one used in Refs. [14, 15].
The remainder of this paper is organized as follows. In Section 2, we outline the formulations employed in this study. Section 3 presents the numerical results and a discussion of their implications. Finally, we provide a summary of our findings in Section 4.
2 Mathematical Formulations
To make the discussion self-contained, in Sect. 2.1, we briefly review the chaotic magnetic field of NSs, while the HT formalism is given in Sect. 2.2. It is important to note that in the mathematical formulations we use $G=c=1$
2.1 Chaotic magnetic field of Neutron Stars
It has been widely understood that the magnetic field could generate anisotropy within NSs [6, 9, 16]. This problem would make the HT formalism for rotational configuration becomes more complex (please see Refs. [1, 17]). For the magnetic field $B$ in the $z$-direction, it is well-known that the stress tensor is written in the form: $diag(\frac{B^{2}}{8\pi};\frac{B^{2}}{8\pi};-\frac{B^{2}}{8\pi})$, being non identical [12]. In Ref. [18], it is argued that the effect of a magnetic field can be described using the concept of pressure only in the case of a small-scale chaotic field. Under this condition, the pressure due to the magnetic field $p_{B}$ is shown to be consistent with field theory. Authors in Ref. [12] agreed to this argument. Now $p_{B}$ reads
$p_{B}=\frac{1}{3}<T^{j}_{j}>=\frac{1}{3}\left(\frac{B^{2}}{8\pi}+\frac{B^{2}}{8\pi}-\frac{B^{2}}{8\pi}\right)=\frac{B^{2}}{24\pi},$ (1)
where $T^{j}_{j}$ denotes the spatial components of the EMT. With this formulation in our hand, we can avoid the anisotropy which is caused by the appearance of magnetic field. The total energy density $\epsilon$ and total pressure $p$ now write
$\epsilon=\epsilon_{m}+\frac{B^{2}}{8\pi},$ (2)
$p=p_{m}+\frac{B^{2}}{24\pi},$ (3)
where the subscript $m$ stands for the matter contribution.
An ansatz of the magnetic field is proposed in Ref. [12], i.e.
$B=B_{0}\left(\frac{\epsilon_{m}}{\epsilon_{0}}\right)^{\gamma}+B_{surf}.$ (4)
Here $B_{0}$ can be interpreted as the magnetic field at the center of the star, $\epsilon_{0}$ is the energy density at the center of the NS with maximum mass when the magnetic field is zero, $\gamma$ is any positive number (in this work we use $\gamma=10^{-3}$)  [^1] Values of $\gamma$ around $10^{-3}$ yield similar results, while setting $\gamma$ to much larger values (e.g., 1, 2, 3) significantly affects the outcome, likely due to changes in energy density affecting the core-crust transition., and $B_{surf}$ is the magnetic field at the surface of the NSs.
2.2 Hartle-Thorne Formalism for Rotating Relativistic Stars
In HT formalism, the metric reads [1, 10, 11, 15]
$ds^{2}$ $=$ $-e^{2\nu}dt^{2}+e^{2\lambda}dr^{2}+r^{2}e^{2\psi}(d\phi-\omega\>dt)^{2}+r^{2}e^{2\mu}d\theta^{2},$ (5)
where $\omega$ is the angular velocity of the local inertial frame, which is proportional to the star’s angular velocity $\Omega$ relative to a distant observer. In this case, $\omega$ and $\Omega$ satisfy $\omega=\Omega-\bar{\omega}$, where $\bar{\omega}$ is the angular velocity of the star relative to the local inertial frame.
Due to rotational perturbation, the exponential functions in Eq. (5) are expanded as the following:
$e^{2\nu}$ $=$ $e^{2\varphi}[1+2(h_{0}+h_{2}P_{2}(\cos\theta))],$ (6)
$e^{2\lambda}$ $=$ $\left[1+\frac{2}{r}\left(m_{0}+m_{2}P_{2}(\cos\theta)\right)\left(1-\frac{2m(r)}{r}\right)^{-1}\right]\left(1-\frac{2m(r)}{r}\right)^{-1}$ (7)
$e^{2\psi}$ $=$ $r^{2}\sin^{2}\theta[1+2(v_{2}-h_{2})P_{2}(\cos\theta)],$ (8)
$e^{2\mu}$ $=$ $r^{2}[1+2(v_{2}-h_{2})P_{2}(\cos\theta)].$ (9)
Here $h_{0}$, $h_{2}$, $m_{0}$, $m_{2}$, and $v_{2}$ are functions of perturbation due to rotation; $P_{2}(\cos\theta)$ is the second order of Legendre polynomial; $e^{2\varphi}$ is a function which is constrained by
$\frac{d\varphi}{dr}=\frac{m+4\pi r^{3}p}{r(r-2m)},$ (10)
where $m$ is the mass function which is constrained by
$\frac{dm}{dr}=4\pi r^{2}\epsilon.$ (11)
The equilibrium equation of the relativistic bodies is given by the Tolman-Oppenheimer-Volkoff (TOV) equation [1, 10, 19], i.e.
$\frac{dp}{dr}=-(\epsilon+p)\frac{d\varphi}{dr}.$ (12)
For a rotating configuration, $\bar{\omega}$ can be calculated by solving the following equation [10]:
$\frac{1}{r^{4}}\frac{d}{dr}\left(r^{4}j\frac{d\bar{\omega}}{dr}\right)+\frac{4}{r}\frac{dj}{dr}=0,$ (13)
where
$j=e^{-\varphi}\left(1-\frac{2m}{r}\right)^{1/2}.$ (14)
The boundary condition at $r=0$ is $\bar{\omega}=\omega_{c}$, where $\omega_{c}$ can be chosen arbitrarily.
It is important to note that Eq. (10), Eq. (11), and Eq. (12) deal with static configuration. Moreover, due to the rotation, the radius of the star is deformed. To obtain the mass and radius corrections of the rotating NS, we have to employ Einstein field equation
$R^{\mu}_{\nu}-\frac{1}{2}\mathscr{R}\delta^{\mu}_{\nu}=8\pi T^{\mu}_{\nu},$ (15)
where $R^{\mu}_{\nu}$ is Ricci tensor, $\mathscr{R}$ is Ricci scalar, $\delta^{\mu}_{\nu}$ is Kronecker delta, and $T^{\mu}_{\nu}$ denotes the EMT.
By calculating the Einstein field equation, for the ($tt$)-component of the $l=0$ order, we can obtain
$\frac{dm_{0}}{dr}$ $=$ $4\pi r^{2}\frac{d\epsilon}{dp}\left(\epsilon+p\right)\mathscr{P}_{0}+\frac{1}{12}j^{2}r^{4}\left(\frac{d\bar{\omega}}{dr}\right)^{2}-\frac{1}{3}r^{3}\left(\frac{dj^{2}}{dr}\right)\bar{\omega}^{2}.$ (16)
From the ($rr$)-component of the $l=0$ order, we can obtain
$\frac{d\mathscr{P}_{0}}{dr}$ $=$ $-\frac{m_{0}\left(1+8\pi r^{2}p\right)}{\left(r-2m\right)^{2}}-\frac{8\pi r^{2}\left(\varepsilon+p\right)}{2\left(r-2m\right)}\mathscr{P}_{0}+\frac{1}{12}\frac{r^{4}j^{2}}{r-2m}\left(\frac{d\bar{\omega}}{dr}\right)^{2}+\frac{1}{3}\left[\frac{d}{dr}\left(\frac{r^{3}j^{2}\bar{\omega}^{2}}{r-2m}\right)\right].$ (17)
Here $\mathscr{P}_{0}$ is originally the 0th-order of expansion of pressure perturbation factor $\mathscr{P}=\mathscr{P}_{0}+\mathscr{P}_{2}P_{2}(\cos\theta)$. The boundary conditions at $r=0$ are $m_{0}(0)=0$ and $\mathscr{P}_{0}(0)=0$.
Mass correction $\delta M$ of the star is given by
$\delta M=m_{0}(R)+\frac{L^{2}}{R^{3}},$ (18)
where $R$ denotes the radius of the star, and $L$ denotes the angular momentum, which satisfies
$L=\frac{R^{3}}{2}\left(\Omega-\bar{\omega}(R)\right).$ (19)
Thus, the total mass reads
$M=M_{STAT}+\delta M,$ (20)
where $M_{STAT}$ is the mass of NSs within static configuration which is obtained by solving Eq. (10), Eq. (11), and Eq. (12) simultneously.
From the ($\theta\theta$), ($\phi$,$\phi$), ($\theta r$), and ($rr$) components of the $l=2$ order of Einstein field equation, we obtain
$\frac{dv_{2}}{dr}=-2\frac{d\varphi}{dr}h_{2}+\left(\frac{1}{r}+\frac{d\varphi}{dr}\right)\left[-\frac{1}{3}r^{3}\frac{dj^{2}}{dr}\bar{\omega}^{2}+\frac{1}{6}j^{2}r^{4}\left(\frac{d\bar{\omega}}{dr}\right)^{2}\right],$ (21)
$\frac{dh_{2}}{dr}$ $=$ $-2\frac{d\varphi}{dr}h_{2}+\frac{\left[8\pi r(\epsilon+p)-\frac{4m}{r^{2}}\right]}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}h_{2}-\frac{4v_{2}}{r(r-2m)\left(2\frac{d\varphi}{dr}\right)}$ (22)
$+\frac{r^{3}j^{2}}{6}\left[\frac{d\varphi}{dr}r-\frac{1}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}\right]\left(\frac{d\bar{\omega}}{dr}\right)^{2}$
$-\frac{r^{2}\bar{\omega}^{2}}{3}\left[\frac{d\varphi}{dr}+\frac{1}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}\right]\frac{dj^{2}}{dr}.$
The boundary conditions at $r=0$ are $v_{2}(0)=0$ and $h_{2}=0$.
Figure 1: Impact of chaotic magnetic field on mass-radius relation of NSs for static configuration.
(a)                                                                (b)
Figure 2: Impact of chaotic magnetic field on mass-radius relation of NSs for (a) polar radius, and (b) equatorial radius.
There exists a formula [1, 10]
$\mathscr{P}_{2}=-h_{2}-\frac{1}{3}r^{2}e^{-2\varphi}\bar{\omega}^{2}.$ (23)
The radius correction of the star is given by
$\delta r=\xi_{0}(R)+\xi_{2}(R)P_{2}(\cos\theta),$ (24)
where
$\xi_{0}=-\mathscr{P}_{0}(\epsilon+p)\left(\frac{dp}{dr}\right)^{-1},$
$\xi_{2}=-\mathscr{P}_{2}(\epsilon+p)\left(\frac{dp}{dr}\right)^{-1}.$ (25)
Finally, we can calculate the radius of the pole $R_{POL}$ and the radius of the equator $R_{EQ}$ [1, 20]
$R_{POL}=R+\xi_{0}+\xi_{2},$ (26)
$R_{EQ}=R+\xi_{0}-\frac{\xi_{2}}{2}.$ (27)
3 Results and Discussion
We numerically solved all differential equations using Euler method. For the value of $\Omega$, we use $\Omega=1000$ s-1. For the initial value of $\bar{\omega}$ (i.e. $\omega_{c}$), we choose 80 s-1 in this work. Please note that this value is much less than the value used in Ref. [21], i.e. $\bar{\omega}$ = 3000 s-1. Moreover, it is also worth noting that $\bar{\omega}(r)$ increases with $r$, which would result in a large $\bar{\omega}$ value at the surface if a large initial $\omega_{c}$ is chosen. On the other side, we have an insight that according to Ref. [22], the HT formalism is accurate to better than 1 per cent even for the fastest millisecond pulsars. However, putting $\bar{\omega}_{c}$ = 3000 s-1 in the calculation is still potential to be dangerous, for the HT formalism is basically an approximation for slowly rotating relativistic stars. So our value of $\bar{\omega}_{c}$ is safer than the one in Ref. [21]. If we compare our initial value of $\bar{\omega}$ to the ones of Ref. [1] and Ref. [14] which chose $\bar{\omega}_{c}=0$, our choice of $\bar{\omega}_{c}$ is more realistic for general rotating NS, since their choice seems to only agree the NSs with very slow rotation. The same value of $\omega_{c}$ and same reasoning of the choice have been highlighted by authors in Ref. [15].
Fig. 1 presents the mass-radius relationship of NSs in a static configuration. The presence of a chaotic magnetic field can reduce the radius of NSs, as indicated by the leftward shift of the curves with increasing magnetic field strength, suggesting that a chaotic magnetic field can contribute to greater compactness of the stars. Compactness, defined as the mass-to-radius ratio, is directly influenced by this field. Notably, the radius reduction is pronounced in the case where $B_{0}=3\times 10^{18}$ G and $B_{surf}=10^{15}$ G, indicating that the radius decrease is non-linear with respect to the increase in the chaotic magnetic field. Furthermore, NSs formed with stronger chaotic magnetic fields exhibit a lower maximum mass compared to those formed with weaker chaotic magnetic fields. Our findings, however, differ from those presented by Lopes et al. [12], who reported that chaotic magnetic fields increase the radius of NSs and allow for the formation of NSs with a higher maximum mass when the chaotic magnetic field is stronger. This discrepancy may be attributed to the different EOS employed in their study. Additionally, it is worth noting that as the magnetic field strength increases, the mass-radius curve begins to resemble that of quark stars. For more detailed discussions on quark stars, please refer to Refs. [23, 24, 25].
Figure 3: Mass-radius relation of the deformed NSs without magnetic field
(a)   $M=0.31\>M_{\odot}$                                              (b)   $M=0.51\>M_{\odot}$ Refer to caption Refer to caption              Refer to caption Refer to caption
(c)   $M=1.18\>M_{\odot}$                                              (d)   $M=1.83\>M_{\odot}$
Figure 4: Illustrations of 3- and 2-dimensional deformation of neutron stars in each particular mass at $B_{0}=3\times 10^{18}$ G and $B_{surf}=1\times 10^{15}$ G.
Fig. 2 shows the polar and equatorial radii, indicating the deformation of rotating NSs. The deformation is less pronounced for the NSs with higher masses but becomes more evident for lower masses. This can be seen in the mass-radius relations of magnetized rotating NSs, where the curves in panel (a) and panel (b) diverge significantly at lower mass values. A comparison with Fig. 3 which depicts rotating NSs without a magnetic field, reveals that deformation still occurs but with only a slight difference between the polar and equatorial radii. This deformation due to rotation is expected and is similar to the Earth’s rotational deformation, where the polar radius is smaller than the equatorial radius [1]. The key takeaway from Fig. 2 and Fig. 3 is that the deformation induced by the magnetic field is much more significant than that caused by rotation.
Additionally, Fig. 2 demonstrates a substantial increase in deformation with increasing magnetic field strength. This contrasts with the findings in Ref. [16], where NSs were also deformed by the presence of a magnetic field, but the increase in deformation with increasing field strength was less pronounced, leading to nearly coincident curves. This difference may stem from differences in the ansatz and core EOS employed in their model.
We present 3- and 2-dimensional illustrations of the NSs in Fig. 4. We take examples from the case of $B_{0}=3\times 10^{18}$ G and $B_{surf}=10^{15}$ G. At lower mass, the star’s shape becomes very oblate. At $M=0.31\>M_{\odot}$, $R_{p}$ = 4.96 km, and $R_{e}$ = 12.51 km; at $M=0.51\>M_{\odot}$, $R_{p}$ = 7.75 km, and $R_{e}$ = 11.99 km; $M=1.18\>M_{\odot}$, $R_{p}$ = 10.67 km, and $R_{e}$ = 11.87 km; and $M=1.83\>M_{\odot}$, $R_{p}$ = 10.48 km, and $R_{e}$ = 10.63 km. It is noteworthy to compare our results with those presented in Ref. [1]. The deformation of rotating NSs in our study is significantly more pronounced than in their findings, which are based on rotating NSs with anisotropic pressure of matter, but without considering the contribution of a magnetic field to the deformation. This fact strengthens our argument that the magnetic field strongly induces the deformation of rotating NSs.
4 Conclusion
In this study, we calculated the mass and radius of NSs under the influence of a chaotic magnetic field, enabling us to obtain the mass-radius relationship for these stars. Our findings indicate that the magnetic field can decrease the radius of NSs. NSs formed with stronger chaotic magnetic fields exhibit a lower maximum mass compared to those formed with weaker chaotic magnetic fields. Additionally, while the chaotic magnetic field can contribute to a reduction in radius, it can simultaneously enhance the compactness and deformation of rotating NSs, with significant deformation occurs at lower mass values.
Acknowledgements
MLP acknowledges the Indonesia Endowment Fund for Education Agency (LPDP) for financial support. Special thanks are extended to Anna Campoy Ordaz for making her code publicly available, which allowed MLP to modify it for use in this work. FPZ and GH would like to thank Kemendikbudristek Republic of Indonesia (through LPPM ITB, Indonesia) for partially financial support.

### v7

@@
-Impact of chaotic magnetic field on mass-radius relation of rotating neutron stars
+Impact of chaotic magnetic field on physical properties of rotating neutron stars
 Abstract
@@
 Apart from being celestial bodies, NSs are normally rotating.
-In this work, we study the impact of a chaotic magnetic field on the mass-radius relation of the rotating NSs.
-We employ an equation of state of NSs with the nuclei in the crust and hyperons in the core.
-We use Hartle-Thorne formalism as an approximation of the rotating NSs.
-For the magnetic field ansatz, we use the one coupled to the energy density.
+In this work, we study the impact of a chaotic magnetic field on the physical properties of the rotating NSs.
+i.e.
+mass, radius, Kepler frequency, and moment of inertia.
+We employ an equation of state (EOS) of NSs with the relativistic mean-field (RMF) model.
+Hartle-Thorne formalism as an approximation of the rotating NSs is utilized.
 We find that the magnetic field can decrease radius of NS.
-NSs formed with stronger chaotic magnetic fields exhibit a lower maximum mass compared to those formed with weaker chaotic magnetic fields.
+NSs formed with magnetic field exhibit a lower maximum mass compared to those without magnetic field.
 In contrast, the increment of the magnetic field can increase the compactness and deformation of rotating NSs.
+The presence of chaotic magnetic field enhances the Kepler frequency of rotating NSs, whereas it simultaneously tend to decreases their moment of inertia.
+The moment of inertia of rotating NSs is consistent with the constraint range obtained from pulsar mass measurements, gravitational wave event data, and X-ray observations of emissions from hotspots on NS surfaces measured by NICER.
 1 Introduction
-It has been widely understood that the evolution of the massive stars whose mass is 8--25 $M_{\odot}$ ended when neutron stars (NSs) formed.
-The corresponding formation is due to the stars’ gravitational collapse during Type-II, Ib, or Ic supernova explosion phenomena [1, 2].
-NSs are dense, compact remnants of evolved stars, with degenerate fermions compressed by strong gravitational fields [3].
-NSs are normally rotating as all celestial objects tend to naturally always rotate due to the law of angular momentum conservation [4].
-Rapidly rotating NSs are observed as pulsars [5].
+It has been widely understood that the evolution of the massive stars whose mass is 8--25 $M_{\odot}$ ends when neutron stars (NSs) form.
+The corresponding formation is due to the stars’ gravitational collapse during Type-II, Ib, or Ic supernova explosion phenomena Pattersons2021 ; Vidaña2018 .
+NSs are compact remnants of evolved stars, with degenerate fermions compressed by strong gravitational fields Reisenegger2001 .
+The stabilizing effect of gravity allows long-time-scale weak interactions (such as electron captures) to reach equilibrium, forming matter that is neutron-rich Graber2017 .
+These stars cover an extensive density range, roughly several times of nuclear density ($n_{0}\sim 10^{14}$ gm/cc), throughout their outer crust to their inner core Ghosh2024 .
+NSs are normally rotating as all celestial objects tend to naturally always rotate due to the law of angular momentum conservation Sivaram2012 .
+When rotating NSs are highly magnetized, they are recognized as pulsars Cohen1972 ; Horvath2004 ; Lyne2013 ; Basu2018 .
 NSs have a wide range of magnetic fields.
-Observations indicate that magnetic fields on NSs are at least in the range of $10^{8-15}$ G.
+Observations indicate that magnetic fields on NSs are at least in the range of $10^{8-15}$ G Reisenegger2001 .
+It is important to note that the direct observation of the magnetic field in NSs’ cores remains unachieved.
+However, theoretical predictions suggest that the strengths of the magnetic fields in NSs’ cores are in the order of $10^{18}$ G to $10^{20}$ G Rizaldy2018b ; Franzon2016 ; Pattersons2024b .
 New classes of pulsars such as Anomalous X-ray pulsars (AXPs) and Soft-Gamma Repeater pulsars (SGRs) have been identified producing vast magnetic fields.
-SGR is associated with remnants of a supernova, which is a young NS [6].
+SGR is associated with remnants of a supernova, which is a young NS Rizaldy2018a .
 Highly magnetized NSs, i.e.
-magnetars, are typically slowly spinning NSs whose emission is thought to be powered by the decay of their large ($\geq 10^{12}$ G) magnetic fields [7].
-Moreover, observations on some AXPs also indicate that AXPs surface magnetic field is around $10^{14}-10^{15}$ G [6].
+magnetars, are typically slowly rotating NSs whose emission is thought to be powered by the decay of their large ($\geq 10^{12}$ G) magnetic fields, much greater than pulsars Desvignes2024 .
+As well as their extreme magnetic fields, magnetars are thought to have core temperature to be well above 109 K Fortin2021 .
+Moreover, observations on some AXPs also indicate that magnetic field of AXPs surface is around $10^{14}-10^{15}$ G Rizaldy2018a .
 Konno et al.
-[8] calculated the deformation of NSs resulting from rotation and the presence of magnetic fields.
+Konno2000 calculated the deformation of NSs resulting from rotation and the presence of magnetic fields.
 In their analysis, a polytropic equation of state (EOS) was employed.
 Their findings indicate that for magnetars, the deformation induced by magnetic fields is dominant, whereas for typical pulsars, the magnetic effect is negligible.
-Mallick & Schramm [9] investigated mass corrections and the deformation of NSs under the influence of magnetic fields, treating the magnetic field as an anisotropy in the energy-momentum tensor (EMT).
-Interestingly, despite focusing on static NSs, their approach was based on the Hartle-Thorne (HT) formalism [10, 11], which is typically applied to rotating relativistic stars.
+Mallick & Schramm Mallick2014 investigated mass corrections and the deformation of NSs under the influence of magnetic fields, treating the magnetic field as an anisotropy in the energy-momentum tensor (EMT).
+Interestingly, despite focusing on static NSs, their approach was based on the Hartle-Thorne (HT) formalism Hartle1967 ; Hartle1968 , which is typically applied to rotating relativistic stars.
 In their work, magnetic perturbations in static configurations were treated analogously to rotational perturbations, following the HT approximation.
 They considered both very stiff and very soft EOS models in their study.
-Lopes & Menezes [12] explored the effects of chaotic magnetic fields on NSs.
+Lopes & Menezes Lopes2015 explored the effects of chaotic magnetic fields on NSs.
 A key advantage of incorporating a chaotic magnetic field lies in the elimination of anisotropy, simplifying the mathematical formulation.
 They also introduced an ansatz where the magnetic field is coupled to the energy density of the NS.
-For their EOS models, they utilized two variations: one including hyperons and one without.
+For their EOS models, they utilized two variations: one including hyperons and one without hyperons.
 However, the rotation of NSs was not considered in their analysis.
-Building on previous studies, we investigate the influence of the chaotic magnetic field on the mass-radius relationship of rotating NSs.
+On the other hand, in Newtonian theory, there exists a maximum rotational frequency for stars composed of perfect fluid, where gravitational and centrifugal forces are in equilibrium at their equators.
+This maximum frequency is proportional to the square root of the ratio of their masses to their radii Wen2007 .
+In the context of general relativity (GR), the angular velocity of NSs is limited by the Kepler frequency, a threshold determined by their mass and radius in static configurations Ordaz2019 ; Lopes2024 .
+Notably, Refs.
+Wen2007 ; Ordaz2019 ; Lopes2024 provide specific calculations of the Kepler frequency.
+Another significant property of NSs is their moment of inertia, which offers insights into their internal structure and EOS.
+Given the link between the EOS and moment of inertia, approximate methods for estimating it in static models, particularly for stiff EOS scenarios, are feasible daSilva2021 .
+Furthermore, measuring the moment of inertia is crucial due to its universal correlation with compactness Rahmansyah2020 .
+Rahmansyah et al.
+Rahmansyah2020 calculated the moment of inertia of anisotropic NS which satisfies the constraints proposed by Landry et al.
+Landry2020 .
+However, the magnetic field is not taken into account in their calculation of the moment of inertia.
+Building on previous studies, we investigate the influence of the chaotic magnetic field on some physical properties of rotating NSs, i.e.
+mass, radius, Kepler frequency, and moment of inertia.
+In particular, for the moment of inertia, we adopt the very slow rotation approximation with angular velocity $\Omega=50$ s-1.
+Under this condition, stellar deformation is neglected, as the stars remain nearly perfectly spherical, with eccentricity values approaching zero.
+This justifies the use of the standard moment of inertia formula for a spherical body.
+Such an approach has also been adopted in Refs.
+Rahmansyah2020 ; Rizaldy2024 .
 The rotational aspects are modeled using the HT formalism, while the magnetic field is described by the Lopes-Menezes ansatz.
-The EOS employed in this work is based on the realistic model proposed by Miyatsu et al.
-[13], which includes nuclei in the crust and hyperons in the core.
+In general, NS has three parts with different compositions and density ranges i.e., outer crust, inner crust and the core Sulaksono2015 .
+These compositions are described by an EOS.
+Here we employ the relativistic mean-field (RMF) EOS with BSP parameter sets Agrawal2012 .
 The algorithm of the numerical simulation follows the one used in Refs.
-[14, 15].
+Ordaz2019 ; Pattersons2024 .
 The remainder of this paper is organized as follows.
@@
 To make the discussion self-contained, in Sect.
-2.1, we briefly review the chaotic magnetic field of NSs, while the HT formalism is given in Sect.
+2.1, we briefly present the nuclear model used in this work; then in Sec 2.2, we shortly review the chaotic magnetic field of NSs; while the HT formalism is given in Sect.
 2.2.
-It is important to note that in the mathematical formulations we use $G=c=1$
+It is important to note that in the mathematical formulations we use $G=c=1$.
+2.1 Nuclear Model
+The Lagrangian density for the RMF model reads Agrawal2012
+$\mathcal{L}$ $=$ $\mathcal{L}_{\mathcal{N}M}+\mathcal{L}_{\sigma}+\mathcal{L}_{\omega}+\mathcal{L}_{\rho}+\mathcal{L}_{\sigma\omega\rho},$ (1)
+where $\mathcal{L}_{\mathcal{N}M}$ is the Lagrangian that describes the interactions of the nucleons through the mesons. It is given by
+$\mathcal{L}_{\mathcal{N}M}$ $=$ $\sum_{J=n,p}\bar{\Psi}_{J}\left[i\gamma^{\mu}\partial_{\mu}-(M-g_{\sigma}\sigma)\right.$ (2)
+$\left.-\left(g_{\omega}\gamma^{\mu}\omega_{\mu}+\frac{1}{2}g_{\rho}\gamma^{\mu}\tau\cdot\rho_{\mu}\right)\right]\Psi_{J}.$
+Note that the sum is taken over the neutrons and protons.
+Here $\tau$ are the isospin matrices.
+Moreover, the Lagrangian that describes the self-interactions for $\sigma$, $\omega$, and $\rho$ mesons can be written as,
+$\mathcal{L}_{\sigma}$ $=$ $\frac{1}{2}(\partial_{\mu}\sigma\partial^{\mu}\sigma-m_{\sigma}^{2}\sigma^{2})-\frac{\kappa_{3}}{6M}g_{\sigma}m_{\sigma}^{2}\sigma^{3}$ (3)
+$-\frac{\kappa_{4}}{24M^{2}}g_{\sigma}^{2}m_{\sigma}^{2}\sigma^{4},$
+$\mathcal{L}_{\omega}$ $=$ $-\frac{1}{4}\omega_{\mu\nu}\omega^{\mu\nu}+\frac{1}{2}m_{\omega}^{2}\omega_{\mu}\omega^{\mu}$ (4)
+$+\frac{1}{24}\zeta_{0}g_{\omega}^{2}(\omega_{\mu}\omega^{\mu})^{2},$
+$\mathcal{L}_{\rho}$ $=$ $-\frac{1}{4}\rho_{\mu\nu}\cdot\rho^{\mu\nu}+\frac{1}{2}m_{\rho}^{2}\rho_{\mu}\cdot\rho^{\mu},$ (5)
+Here, $\omega_{\mu\nu}$ and $\rho^{\mu\nu}$ are the field tensors corresponding to the $\omega$ and $\rho$ mesons, respectively.
+They are defined as $\omega_{\mu\nu}=\partial_{\mu}\omega_{\nu}-\partial_{\nu}\omega_{\mu}$ and $\rho^{\mu\nu}=\partial^{\mu}\rho^{\nu}-\partial^{\nu}\rho^{\mu}$.
+Furthermore, the cross interactions of $\sigma$, $\omega$, and $\rho$ mesons are given by $\mathcal{L}_{\sigma\omega\rho}$, i.e.
+$\mathcal{L}_{\sigma\omega\rho}$ $=$ $\frac{\eta_{1}}{2M}g_{\sigma}m_{\omega}^{2}\sigma\omega_{\mu}\omega^{\mu}+\frac{\eta_{2}}{4M^{2}}g_{\sigma}^{2}m_{\omega}^{2}\sigma^{2}\omega_{\mu}\omega^{\mu}$ (6)
+$+\frac{\eta_{\rho}}{2M}g_{\sigma}m_{\rho}^{2}\sigma\,\rho_{\mu}\cdot\rho^{\mu}$
+$+\frac{\eta_{1\rho}}{4M^{2}}g_{\sigma}^{2}m_{\rho}^{2}\sigma^{2}\,\rho_{\mu}\cdot\rho^{\mu}$
+$+\frac{\eta_{2\rho}}{4M^{2}}g_{\omega}^{2}m_{\rho}^{2}\omega_{\mu}\omega^{\mu}\,\rho_{\mu}\cdot\rho^{\mu}.$
+Table 1: BSP parameter sets Agrawal2012 for the RMF model.
+Value
+$g_{\sigma}/4\pi$ 0.8764
+$g_{\omega}/4\pi$ 1.1481
+$g_{\rho}/4\pi$ 1.0508
+$\kappa_{3}$ 1.0681
+$\kappa_{4}$ 14.9857
+$\eta_{1}$ 0.0872
+$\eta_{2}$ 3.1265
+$\eta_{\rho}$ 0.0
+$\eta_{1\rho}$ 0.0
+$\eta_{2\rho}$ 53.7642
+$\zeta_{0}$ 0.0
+$m_{\sigma}/M$ 0.5383
+$m_{\omega}/M$ 0.8333
+$m_{\rho}/M$ 0.8200
+The values of the free parameters of the Lagrangian for this model are listed in Table 1, while the predicted properties at saturation density—namely the binding energy per baryon ($B/A$), incompressibility ($K$), symmetry energy ($S_{0}$), the slope of the symmetry energy ($L$), and saturation density ($n_{0}$) itself—are summarized in Table 2.
+The constraints for the predicted values at saturation density are taken from Refs.
+Lopes2024 ; Rahmansyah2020 ; Dutra2014 .
+Table 2: Predictions for the RMF EOS with BSP parameter sets at saturation density. the constraints are taken from Refs. Lopes2024 ; Rahmansyah2020 ; Dutra2014 .
+Physical Quantities Constraints This Model
+$n_{0}$ (fm-3) 0.148–0.170 0.149
+$K$ (MeV) 220–260 230.97
+$S_{0}$ (MeV) 28.5–34.9 28.86
+$B/A$ (MeV) 15.8–16.5 15.90
+$L$ (MeV) 44–116 50.00
+For the crust of NSs, the one in Refs Mart2013 ; Suparti2017 is used.
+At the core-crust transition, the pressures of the core and the inner crust are matched.
+The core-crust transition occurs at the transition nuclear density $n_{t}=$ 0.087 fm-3 and transition pressure $p_{t}=$ 0.29 MeV fm-3 Rahmansyah2020 .
+It is important to note that the effect of the magnetic field on the EOS is not taken into account, as our focus is solely on investigating the impact of the chaotic magnetic field on general relativistic aspects of NSs and their implication on the NSs’ physical properties.
 2.2 Chaotic Magnetic Field of Neutron Stars
-It has been widely understood that the magnetic field could generate anisotropy within NSs [6, 9, 16].
+Generally, chaotically tangled magnetic field lines are thought to be present throughout astrophysical plasmas: planets, stars, accretion discs, galaxies, clusters of galaxies, and the intergalactic medium Kumar2017 .
+In nature, chaotic magnetic fields could be generated by asymmetric current configuration Dasgupta2014 .
+It has been widely understood that the magnetic field could generate anisotropy within NSs Rizaldy2018b ; Rizaldy2018a ; Mallick2014 .
 This problem would make the HT formalism for rotational configuration becomes more complex (please see Refs.
-[1, 17]).
-For the magnetic field $B$ in the $z$-direction, it is well-known that the stress tensor is written in the form: $diag(\frac{B^{2}}{8\pi};\frac{B^{2}}{8\pi};-\frac{B^{2}}{8\pi})$, being non identical [12].
+Pattersons2021 ; Posada2024 ; Becerra2024 ).
+For the magnetic field $B$ in the $z$-direction, it is well-known that the stress tensor is written in the form: $diag(\frac{B^{2}}{8\pi};\frac{B^{2}}{8\pi};-\frac{B^{2}}{8\pi})$, being non identical Lopes2015 .
 In Ref.
-[18], it is argued that the effect of a magnetic field can be described using the concept of pressure only in the case of a small-scale chaotic field.
+Zeldovich , it is argued that the effect of a magnetic field can be described using the concept of pressure only in the case of a small-scale chaotic field.
 Under this condition, the pressure due to the magnetic field $p_{B}$ is shown to be consistent with field theory.
-Authors in Ref.
-[12] agreed to this argument.
+Ref.
+Lopes2015 agreed to this argument.
 Now $p_{B}$ reads
@@
 where the subscript $m$ stands for the matter contribution.
-An ansatz of the magnetic field is proposed in Ref. [12], i.e.
+An ansatz of the magnetic field is proposed in Ref. Lopes2015 , i.e.
 $B=B_{0}\left(\frac{\epsilon_{m}}{\epsilon_{0}}\right)^{\gamma}+B_{surf}.$ (10)
-Here $B_{0}$ can be interpreted as the magnetic field at the center of the star, $\epsilon_{0}$ is the energy density at the center of the NS with maximum mass when the magnetic field is zero, $\gamma$ is any positive number (in this work we use $\gamma=10^{-3}$)  [^1] Values of $\gamma$ around $10^{-3}$ yield similar results, while setting $\gamma$ to much larger values (e.g., 1, 2, 3) significantly affects the outcome, likely due to changes in energy density affecting the core-crust transition., and $B_{surf}$ is the magnetic field at the surface of the NSs.
+Here $B_{0}$ can be interpreted as the magnetic field at the center of the star, $\epsilon_{0}$ is the energy density at the center of the NS with maximum mass when the magnetic field is zero, $\gamma$ is any positive number (in this work we use $\gamma=$ 3, 5, and 6), and $B_{surf}$ is the magnetic field at the surface of the NSs.
 2.3 Hartle-Thorne Formalism for Rotating Relativistic Stars
-In HT formalism, the metric reads [1, 10, 11, 15]
-$ds^{2}$ $=$ $-e^{2\nu}dt^{2}+e^{2\lambda}dr^{2}+r^{2}e^{2\psi}(d\phi-\omega\>dt)^{2}+r^{2}e^{2\mu}d\theta^{2},$ (5)
+The details of the HT formalism can be found in the original papers Hartle1967 ; Hartle1968 .
+Other interesting additional discussions on the formalism can be referred to Refs.
+Pattersons2021 ; Pattersons2024 ; Posada2024 ; Becerra2024 .
+Here we present only a little discussion and the necessary equations to calculate some physical properties, i.e.
+mass, radius, Kepler frequency, and moment of inertia.
+This is especially useful to help the readers who are not familiar with the HT formalism.
+It has to be noted that based on Eq.
+(8) and Eq.
+(9), it can be seen that the total energy density and total pressure arising from the matter and the presence of a magnetic field can be redefined by introducing $\epsilon$ and $p$.
+Therefore, in this subsection, the definitions of $\epsilon$ and $p$ also encompass the total contributions from both matter and the magnetic field.
+Consequently, the formulation of HT presented in this subsection is indirectly modified by the presence of the magnetic field.
+Although this modification is not explicitly apparent, the contribution of the magnetic field is embedded within $\epsilon$ and $p$.
+In HT formalism, the metric reads Pattersons2021 ; Hartle1967 ; Hartle1968 ; Pattersons2024
+$ds^{2}$ $=$ $-e^{2\nu}dt^{2}+e^{2\lambda}dr^{2}+e^{2\psi}(d\phi-\omega\>dt)^{2}$ (11)
+$+e^{2\mu}d\theta^{2},$
 where $\omega$ is the angular velocity of the local inertial frame, which is proportional to the star’s angular velocity $\Omega$ relative to a distant observer.
@@
 $e^{2\nu}$ $=$ $e^{2\varphi}[1+2(h_{0}+h_{2}P_{2}(\cos\theta))],$ (12)
-$e^{2\lambda}$ $=$ $\left[1+\frac{2}{r}\left(m_{0}+m_{2}P_{2}(\cos\theta)\right)\left(1-\frac{2m(r)}{r}\right)^{-1}\right]\left(1-\frac{2m(r)}{r}\right)^{-1}$ (7)
+$e^{2\lambda}$ $=$ $\left[1+\frac{2}{r}\left(m_{0}+m_{2}P_{2}(\cos\theta)\right)\left(1-\frac{2m(r)}{r}\right)^{-1}\right]$ (13)
+×$\times\left(1-\frac{2m(r)}{r}\right)^{-1}$
 $e^{2\psi}$ $=$ $r^{2}\sin^{2}\theta[1+2(v_{2}-h_{2})P_{2}(\cos\theta)],$ (14)
@@
 $\frac{dm}{dr}=4\pi r^{2}\epsilon.$ (17)
-The equilibrium equation of the relativistic bodies is given by the Tolman-Oppenheimer-Volkoff (TOV) equation [1, 10, 19], i.e.
+The equilibrium equation of the relativistic bodies is given by the Tolman-Oppenheimer-Volkoff (TOV) equation Pattersons2021 ; Hartle1967 ; Oppenheimer1939 , i.e.
 $\frac{dp}{dr}=-(\epsilon+p)\frac{d\varphi}{dr}.$ (18)
-For a rotating configuration, $\bar{\omega}$ can be calculated by solving the following equation [10]:
+For a rotating configuration, $\bar{\omega}$ can be calculated by solving the following equation Hartle1967 :
 $\frac{1}{r^{4}}\frac{d}{dr}\left(r^{4}j\frac{d\bar{\omega}}{dr}\right)+\frac{4}{r}\frac{dj}{dr}=0,$ (19)
@@
 The boundary condition at $r=0$ is $\bar{\omega}=\omega_{c}$, where $\omega_{c}$ can be chosen arbitrarily.
-It is important to note that Eq.
-(10), Eq.
-(11), and Eq.
-(12) deal with static configuration.
+Figure 1: Mass-radius relation of NSs for static configuration with different parameter sets of chaotic magnetic field strengths.
+(a)                                                   (b)
+Refer to caption Refer to caption
+                    (c)                                                   (d)
+Refer to caption Refer to caption
+                    (e)                                                   (f)
+Figure 2: Mass-radius relation of NSs for (a) polar radius at $\Omega=$ 1000 s-1, (b) equatorial radius at $\Omega=$ 1000 s-1, (c) polar radius at $\Omega=$ 2000 s-1, (d) equatorial radius at $\Omega=$ 2000 s-1, (e) polar radius at $\Omega=$ 3000 s-1, and (f) equatorial radius at $\Omega=$ 3000 s-1.
+It is important to note that Eqs.
+(16)–(18) deal with static configuration.
 Moreover, due to the rotation, the radius of the star is deformed.
@@
 where $R^{\mu}_{\nu}$ is Ricci tensor, $\mathscr{R}$ is Ricci scalar, $\delta^{\mu}_{\nu}$ is Kronecker delta, and $T^{\mu}_{\nu}$ denotes the EMT.
-By calculating the Einstein field equation, for the ($tt$)-component of the $l=0$ order, we can obtain
-$\frac{dm_{0}}{dr}$ $=$ $4\pi r^{2}\frac{d\epsilon}{dp}\left(\epsilon+p\right)\mathscr{P}_{0}+\frac{1}{12}j^{2}r^{4}\left(\frac{d\bar{\omega}}{dr}\right)^{2}-\frac{1}{3}r^{3}\left(\frac{dj^{2}}{dr}\right)\bar{\omega}^{2}.$ (16)
-From the ($rr$)-component of the $l=0$ order, we can obtain
-$\frac{d\mathscr{P}_{0}}{dr}$ $=$ $-\frac{m_{0}\left(1+8\pi r^{2}p\right)}{\left(r-2m\right)^{2}}-\frac{8\pi r^{2}\left(\varepsilon+p\right)}{2\left(r-2m\right)}\mathscr{P}_{0}+\frac{1}{12}\frac{r^{4}j^{2}}{r-2m}\left(\frac{d\bar{\omega}}{dr}\right)^{2}+\frac{1}{3}\left[\frac{d}{dr}\left(\frac{r^{3}j^{2}\bar{\omega}^{2}}{r-2m}\right)\right].$ (17)
+By calculating the Einstein field equation, for the ($tt$) component of the $l=0$ order, we can obtain
+$\frac{dm_{0}}{dr}$ $=$ $4\pi r^{2}\frac{d\epsilon}{dp}\left(\epsilon+p\right)\mathscr{P}_{0}+\frac{1}{12}j^{2}r^{4}\left(\frac{d\bar{\omega}}{dr}\right)^{2}$ (22)
+$-\frac{1}{3}r^{3}\left(\frac{dj^{2}}{dr}\right)\bar{\omega}^{2}.$
+From the ($rr$) component of the $l=0$ order, we can obtain
+$\frac{d\mathscr{P}_{0}}{dr}$ $=$ $-\frac{m_{0}\left(1+8\pi r^{2}p\right)}{\left(r-2m\right)^{2}}-\frac{8\pi r^{2}\left(\varepsilon+p\right)}{2\left(r-2m\right)}\mathscr{P}_{0}+\frac{1}{12}$ (23)
+×$\times\frac{r^{4}j^{2}}{r-2m}\left(\frac{d\bar{\omega}}{dr}\right)^{2}+\frac{1}{3}\left[\frac{d}{dr}\left(\frac{r^{3}j^{2}\bar{\omega}^{2}}{r-2m}\right)\right].$
 Here $\mathscr{P}_{0}$ is originally the 0th-order of expansion of pressure perturbation factor $\mathscr{P}=\mathscr{P}_{0}+\mathscr{P}_{2}P_{2}(\cos\theta)$.
@@
 $\delta M=m_{0}(R)+\frac{L^{2}}{R^{3}},$ (24)
-where $R$ denotes the radius of the star, and $L$ denotes the angular momentum, which satisfies
-$L=\frac{R^{3}}{2}\left(\Omega-\bar{\omega}(R)\right).$ (19)
+where $R$ denotes the radius of the star, and $L$ denotes the angular momentum. In Ref. Hartle1968 , there is a relation
+${\Omega=\bar{\omega}(R)+\frac{2L}{R^{3}},}$ (25)
+so we can obtain
+$L=\frac{R^{3}}{2}(\Omega-\bar{\omega}(R)).$ (26)
+Alternatively, angular momentum can be calculated as Glendenning1998
+[…]
 Thus, the total mass reads
-$M=M_{STAT}+\delta M,$ (20)
-where $M_{STAT}$ is the mass of NSs within static configuration which is obtained by solving Eq. (10), Eq. (11), and Eq. (12) simultneously.
-From the ($\theta\theta$), ($\phi$,$\phi$), ($\theta r$), and ($rr$) components of the $l=2$ order of Einstein field equation, we obtain
-$\frac{dv_{2}}{dr}=-2\frac{d\varphi}{dr}h_{2}+\left(\frac{1}{r}+\frac{d\varphi}{dr}\right)\left[-\frac{1}{3}r^{3}\frac{dj^{2}}{dr}\bar{\omega}^{2}+\frac{1}{6}j^{2}r^{4}\left(\frac{d\bar{\omega}}{dr}\right)^{2}\right],$ (21)
-$\frac{dh_{2}}{dr}$ $=$ $-2\frac{d\varphi}{dr}h_{2}+\frac{\left[8\pi r(\epsilon+p)-\frac{4m}{r^{2}}\right]}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}h_{2}-\frac{4v_{2}}{r(r-2m)\left(2\frac{d\varphi}{dr}\right)}$ (22)
-$+\frac{r^{3}j^{2}}{6}\left[\frac{d\varphi}{dr}r-\frac{1}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}\right]\left(\frac{d\bar{\omega}}{dr}\right)^{2}$
+$M=M_{0}+\delta M,$ (28)
+where $M_{0}$ is the mass of NSs within static configuration which is obtained by solving Eqs. (16)–(18) simultaneously.
+From the ($\theta\theta$), ($\phi\phi$), ($\theta r$), and ($rr$) components of the $l=2$ order of Einstein field equation, we obtain
+$\frac{dv_{2}}{dr}$ $=$ $-2\frac{d\varphi}{dr}h_{2}+\left(\frac{1}{r}+\frac{d\varphi}{dr}\right)$ (29)
+×$\times\left[-\frac{1}{3}r^{3}\frac{dj^{2}}{dr}\bar{\omega}^{2}+\frac{1}{6}j^{2}r^{4}\left(\frac{d\bar{\omega}}{dr}\right)^{2}\right],$
+$\frac{dh_{2}}{dr}$ $=$ $-2\frac{d\varphi}{dr}h_{2}+\frac{\left[8\pi r(\epsilon+p)-\frac{4m}{r^{2}}\right]}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}h_{2}$ (30)
+$-\frac{4v_{2}}{r(r-2m)}\left(2\frac{d\varphi}{dr}\right)^{-1}+\frac{r^{3}j^{2}}{6}$
+×$\times\left[\frac{d\varphi}{dr}r-\frac{1}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}\right]\left(\frac{d\bar{\omega}}{dr}\right)^{2}$
 $-\frac{r^{2}\bar{\omega}^{2}}{3}\left[\frac{d\varphi}{dr}+\frac{1}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}\right]\frac{dj^{2}}{dr}.$
 The boundary conditions at $r=0$ are $v_{2}(0)=0$ and $h_{2}=0$.
-Figure 1: Impact of chaotic magnetic field on mass-radius relation of NSs for static configuration.
-(a)                                                                (b)
-Figure 2: Impact of chaotic magnetic field on mass-radius relation of NSs for (a) polar radius, and (b) equatorial radius.
-There exists a formula [1, 10]
+There exists a formula Pattersons2021 ; Hartle1967
 $\mathscr{P}_{2}=-h_{2}-\frac{1}{3}r^{2}e^{-2\varphi}\bar{\omega}^{2}.$ (31)
@@
 $\xi_{2}=-\mathscr{P}_{2}(\epsilon+p)\left(\frac{dp}{dr}\right)^{-1}.$ (33)
-Finally, we can calculate the radius of the pole $R_{POL}$ and the radius of the equator $R_{EQ}$ [1, 20]
+Finally, we can calculate the radius of the pole $R_{POL}$ and the radius of the equator $R_{EQ}$ Pattersons2021 ; Lopes2024
 $R_{POL}=R+\xi_{0}+\xi_{2},$ (34)
 $R_{EQ}=R+\xi_{0}-\frac{\xi_{2}}{2}.$ (35)
+This allows us to calculate the eccentricity
+$e=\sqrt{1-\frac{R_{POL}^{2}}{R_{EQ}^{2}}}.$ (36)
+Furthermore, in the original papers of HT formalism Hartle1967 ; Hartle1968 , the maximum angular velocity of a rotating body within Newtonian gravity is mentioned, i.e.
+$\Omega_{max}=\sqrt{M_{0}/R^{3}}.$ (37)
+However, in the context of GR, the maximum angular velocity of NSs (which is called as Kepler frequency) is approximated by Lopes2024
+$\Omega_{K}=0.65\sqrt{M_{0}/R^{3}}.$ (38)
+Moreover, another important physical quantity to consider is the moment of inertia.
+With the angular momentum $L$ and the angular velocity relative to the distant observers $\Omega$ in our hands, we can determine the moment of inertia
+$I=\frac{L}{\Omega}.$ (39)
 3 Results and Discussion
+Table 3: Parameter sets of the magnetic fields
+$\gamma$
+Set A $B=0$
+Set B 3
+Set C 5
+Set D 6
 We numerically solved all differential equations using Euler method.
-For the value of $\Omega$, we use $\Omega=1000$ s-1.
-For the initial value of $\bar{\omega}$ (i.e.
-$\omega_{c}$), we choose 80 s-1 in this work.
-Please note that this value is much less than the value used in Ref.
-[21], i.e.
-$\bar{\omega}$ = 3000 s-1.
+For the value of $\Omega$, we consider $\Omega=1000$ s-1, $\Omega=2000$ s-1, and $\Omega=3000$ s-1.
+However, except for the calculation of the moment of inertia, we use $\Omega=50$ s-1, as we only employ the very slow rotation approximation for this purpose.
+For the initial value of $\bar{\omega}$ (i.e., $\omega_{c}$), we adopt $\omega_{c}=80$ s-1 in all cases throughout this work, except in the calculation of the moment of inertia, where we set $\omega_{c}$.
+It is worth noting that the value $\omega_{c}=0$ used in our calculations is significantly smaller than the value used in Ref.
+Wen , where $\bar{\omega}=3000$ s-1.
 Moreover, it is also worth noting that $\bar{\omega}(r)$ increases with $r$, which would result in a large $\bar{\omega}$ value at the surface if a large initial $\omega_{c}$ is chosen.
-On the other side, we have an insight that according to Ref.
-[22], the HT formalism is accurate to better than 1 per cent even for the fastest millisecond pulsars.
+On the other side, we have insight that according to Ref.
+Ber , the HT formalism is accurate to better than 1 per cent even for the fastest millisecond pulsars.
 However, putting $\bar{\omega}_{c}$ = 3000 s-1 in the calculation is still potential to be dangerous, for the HT formalism is basically an approximation for slowly rotating relativistic stars.
 So our value of $\bar{\omega}_{c}$ is safer than the one in Ref.
-[21].
-If we compare our initial value of $\bar{\omega}$ to the ones of Ref.
-[1] and Ref.
-[14] which chose $\bar{\omega}_{c}=0$, our choice of $\bar{\omega}_{c}$ is more realistic for general rotating NS, since their choice seems to only agree the NSs with very slow rotation.
-The same value of $\omega_{c}$ and same reasoning of the choice have been highlighted by authors in Ref.
-[15].
+Wen .
+Table 4: Macroscopic properties of neutron stars under different magnetic field configurations .
+The quantities are the maximum mass $M^{\text{max}}$ (in $M_{\odot}$), average radius at maximum mass $\bar{R}^{\text{max}}$ (in km), eccentricity at maximum max $e^{\textbf{max}}$, and central energy density at maximum mass $\epsilon^{\text{max}}_{c}$ (in MeV fm-3).
+Additional columns show the corresponding values at $1.4M_{\odot}$.
+$M^{\text{max}}$/$M_{\odot}$ $\bar{R}^{\text{max}}$ (km) $e^{\textbf{max}}$ $\epsilon_{c}^{\text{max}}$ (MeV fm-3) $\bar{R}^{1.4}$ (km) $e^{\text{1.4}}$ $\epsilon^{\text{1.4}}_{c}$ (MeV fm-3)
+Set A at $\Omega=0$ 2.02575 10.74600 0 1309.60 11.85700 0 436.09
+Set A at $\Omega=1000$ s-1 2.02686 10.73685 0.12437 1309.60 11.75275 0.34570 436.09
+Set A at $\Omega=2000$ s-1 2.03020 10.70925 0.24618 1309.60 11.44003 0.63948 436.09
+Set A at $\Omega=3000$ s-1 2.03576 10.66335 0.36303 1309.60 10.91882 0.84868 436.09
+Set B at $\Omega=0$ 1.91962 10.09100 0 1487.32 10.98300 0 525.93
+Set B at $\Omega=1000$ s-1 1.92058 10.07290 0.15953 1487.32 10.83280 0.40704 525.93
+Set B at $\Omega=2000$ s-1 1.92345 10.01860 0.31385 1487.32 10.38220 0.73110 525.93
+Set B at $\Omega=3000$ s-1 1.92824 9.92813 0.45821 1487.32 9.63110 0.92837 525.93
+Set C at $\Omega=0$ 1.91962 10.09100 0 1487.32 10.98300 0 525.93
+Set C at $\Omega=1000$ s-1 1.92058 10.07290 0.15953 1487.32 10.83280 0.40703 525.93
+Set C at $\Omega=2000$ s-1 1.92345 10.01860 0.31384 1487.32 10.38220 0.73109 525.93
+Set C at $\Omega=3000$ s-1 1.92824 9.92816 0.45820 1487.32 9.63120 0.92836 525.93
+Set D at $\Omega=0$ 1.91962 10.09100 0 1487.33 10.98300 0 525.93
+Set D at $\Omega=1000$ s-1 1.92058 10.07290 0.15953 1487.33 10.83280 0.40703 525.93
+Set D at $\Omega=2000$ s-1 1.92345 10.01864 0.31384 1487.33 10.38218 0.73109 525.93
+Set D at $\Omega=3000$ s-1 1.92824 9.92816 0.45820 1487.33 9.63115 0.92837 525.93
+If we compare our initial value of $\bar{\omega}$—excluding the one used for the moment of inertia calculation—to those in Ref.
+Pattersons2021 and Ref.
+Ordaz2019 , which adopted $\bar{\omega}_{c}=0$, our choice of $\bar{\omega}_{c}=80$ s-1 appears to be more realistic for modeling general rotating NSs.
+In contrast, their assumption of $\bar{\omega}_{c}=0$ is more appropriate only for very slowly rotating neutron stars.
+A similar choice of $\omega_{c}$ and the reasoning behind it have also been emphasized by the authors in Ref.
+Pattersons2024 .
+We set the limitation of $\Omega$ to be 3000 s-1 in our analysis.
+Moreover, the parameter sets for the magnetic field employed in this study are summarized in Table 3.
+We consider both scenarios: without and with the presence of a magnetic field.
+For the cases where the magnetic field is present, we vary the parameter $\gamma$ with values of 3, 5, and 6.
+It is also important to note that although $\gamma$ is formally a free parameter and can, in principle, take any positive value, choosing a very small value (e.g., of the order of 10-3) is problematic.
+In such cases, the magnetic field becomes strongly dominated by the constant component $B_{0}$, making it effectively resemble a uniform magnetic field, which is unphysical in the context of our model.
+Figure 3: Mass-radius relation of the deformed NSs without magnetic field.
+(a)   $M=0.9\>M_{\odot}$                                              (b)   $M=1.4\>M_{\odot}$ Refer to caption Refer to caption
+(c)   $M=1.7\>M_{\odot}$                                              (d)   $M=1.9\>M_{\odot}$
+Figure 4: Illustrations of 3- and 2-dimensional deformation of neutron stars in each particular mass when $\gamma=3$ and $\Omega=$ 1000 s-1.
+It has to be noted that the impact of varying the parameter $\gamma$ is not significant.
+As shown in Table 4, we provide examples of how changes in $\gamma$ affect several key quantities, namely the maximum mass $M^{\text{max}}$, the average radius at maximum mass $\bar{R}^{\text{max}}$, and the average radius at the canonical mass $\bar{R}^{1.4}$.
+We adopt the definition of average radius $\bar{R}$ from Ref.
+Pattersons2021 , given by $\bar{R}=\frac{R_{\text{POL}}+R_{\text{EQ}}}{2}$.
+In this context, investigating the canonical mass is particularly relevant, as observational data suggest that NSs predominantly reside within a relatively narrow mass range around $1.4M_{\odot}$ Kaper2006 .
+As shown in Table 4, the variation of $\gamma$ affects the average radius at both the maximum mass and the canonical mass only at the order of $10^{-5}$, and even then, only for a subset of the cases.
+Most of the values remain unchanged.
+In fact, the maximum mass values are completely unaffected by variations in $\gamma$, indicating that its influence on the overall stellar structure is minimal within the explored range.
 Fig.
 1 presents the mass-radius relationship of NSs in a static configuration.
-The presence of a chaotic magnetic field can reduce the radius of NSs, as indicated by the leftward shift of the curves with increasing magnetic field strength, suggesting that a chaotic magnetic field can contribute to greater compactness of the stars.
+This relation purely comes from TOV equation.
+From the relation, we can see that the presence of a chaotic magnetic field can reduce the radius of NSs, as indicated by the leftward shift of the curves with the appearance of magnetic field strength, suggesting that a chaotic magnetic field can contribute to greater compactness of the stars.
 Compactness, defined as the mass-to-radius ratio, is directly influenced by this field.
-Notably, the radius reduction is pronounced in the case where $B_{0}=3\times 10^{18}$ G and $B_{surf}=10^{15}$ G, indicating that the radius decrease is non-linear with respect to the increase in the chaotic magnetic field.
-Furthermore, NSs formed with stronger chaotic magnetic fields exhibit a lower maximum mass compared to those formed with weaker chaotic magnetic fields.
-Our findings, however, differ from those presented by Lopes et al.
-[12], who reported that chaotic magnetic fields increase the radius of NSs and allow for the formation of NSs with a higher maximum mass when the chaotic magnetic field is stronger.
-This discrepancy may be attributed to the different EOS employed in their study.
-Additionally, it is worth noting that as the magnetic field strength increases, the mass-radius curve begins to resemble that of quark stars.
-For more detailed discussions on quark stars, please refer to Refs.
-[23, 24, 25].
-Figure 3: Mass-radius relation of the deformed NSs without magnetic field
-(a)   $M=0.31\>M_{\odot}$                                              (b)   $M=0.51\>M_{\odot}$ Refer to caption Refer to caption              Refer to caption Refer to caption
-(c)   $M=1.18\>M_{\odot}$                                              (d)   $M=1.83\>M_{\odot}$
-Figure 4: Illustrations of 3- and 2-dimensional deformation of neutron stars in each particular mass at $B_{0}=3\times 10^{18}$ G and $B_{surf}=1\times 10^{15}$ G.
+Notably, the maximum mass of the NSs also decreases when the magnetic field appears.
+Furthermore, when comparing our results with previous works that also employ chaotic magnetic fields using the same ansatz—such as Refs.
+Lopes2015 ; Wu2017 ; Lopes2020 —a notable difference arises.
+While those studies report that the presence of chaotic magnetic fields allows for NSs with higher maximum masses, our results show a reduction in the maximum mass when magnetic fields are included.
+This discrepancy may be attributed to the fact that we do not include the magnetic field in the EOS, whereas the aforementioned studies do.
+As a result, their EOS becomes intrinsically stiffer in the presence of magnetic fields, which naturally leads to higher maximum masses.
+Nevertheless, in terms of the shape of the mass-radius relation, particularly when varying $\gamma$ within the range of 3 to 6, our results are still in qualitative agreement with the aforementioned studies.
+For example, while Ref.
+Wu2017 shows more noticeable deviations between curves with different $\gamma$ values, the results in Refs.
+Lopes2015 and Lopes2020 display behavior more similar to ours, where the differences are relatively minor and the curves seem to coincide.
+These variations may also stem from differences in the underlying EOS models, as it is well known that different EOS models can lead to significantly different outcomes in NS structure calculations Rather2021 .
+In addition to the nuclear matter constraints, we also consider astrophysical constraints in this work, namely those related to PSR J0740+6620 and the radius at the canonical mass.
+PSR J0740+6620 has a mass of $2.08\pm 0.07$ $M_{\odot}$ Miller2021 , with a radius of $12.39^{+0.98}_{-1.30}$ km Riley2021 .
+As shown in Fig.
+1, the maximum mass of NSs without magnetic fields satisfies the lower bound of the mass constraint, while the radius constraint is not fulfilled.
+NSs with magnetic fields fail to meet both constraints.
+Regarding the radius at canonical mass, neutron stars without magnetic fields satisfy the constraint, whereas those with magnetic fields do not.
+The exploration of an EOS that satisfies all nuclear and astrophysical constraints is left for future work.
+Figure 5: Eccentricity of NSs at (a) $\Omega=$ 1000 s-1, (b) $\Omega=$ 2000 s-1, and (c) $\Omega=$ 3000 s-1 as a function of mass.
+Figure 6: Kepler frequency of NSs as a function of mass.
+(a)                                                                          (b)
+Figure 7: (a) Moment of inertia of NSs when $\Omega=$ 50 s-1, (b) Eccentricity of the NSs when $\Omega=$ 50 s-1.
+Here we use a constraint range for moment of inertia presented by Landry et al.
+Landry2020 .
 Fig.
@@
 The deformation is less pronounced for the NSs with higher masses but becomes more evident for lower masses.
-This can be seen in the mass-radius relations of magnetized rotating NSs, where the curves in panel (a) and panel (b) diverge significantly at lower mass values.
+This can be seen in the mass-radius relations of magnetized rotating NSs, where the curves in $R_{POL}$ panels and $R_{EQ}$ panels diverge significantly at lower mass values.
 A comparison with Fig.
-3 which depicts rotating NSs without a magnetic field, reveals that deformation still occurs but with only a slight difference between the polar and equatorial radii.
-This deformation due to rotation is expected and is similar to the Earth’s rotational deformation, where the polar radius is smaller than the equatorial radius [1].
+3 which depicts rotating NSs without a magnetic field, reveals that deformation still occurs but with only slighter differences between the polar and equatorial radii.
+This deformation due to rotation is expected and is similar to the Earth’s rotational deformation, where the polar radius is smaller than the equatorial radius Pattersons2021 .
 The key takeaway from Fig.
@@
 3 is that the deformation induced by the magnetic field is much more significant than that caused by rotation.
+It is worth noting that even though the magnetic field strengthens the deformation of the stars, the magnetic field in the static case shown by Fig.
+1 does not cause the deformation, because the deformation in this context comes from quadrupole order of HT formalism.
+Since the magnetic field is embedded within the energy density and pressure which exist in the perturbative terms of HT formalism, so the magnetic field could impact on the deformation of the rotating NSs.
+It is important to emphasize that the rotational aspect also plays an important role to deform the NSs.
+As shown in Fig.
+2 and Fig.
+3, higher values of $\Omega$ lead to a greater separation between the $R_{\text{POL}}$ and $R_{\text{EQ}}$ curves.
 Additionally, Fig.
-2 demonstrates a substantial increase in deformation with increasing magnetic field strength.
-This contrasts with the findings in Ref.
-[16], where NSs were also deformed by the presence of a magnetic field, but the increase in deformation with increasing field strength was less pronounced, leading to nearly coincident curves.
-This difference may stem from differences in the ansatz and core EOS employed in their model.
+2 demonstrates a substantial increase in deformation with the appearance of the magnetic field.
+This agrees with the findings in Ref.
+Rizaldy2018b , where NSs were also deformed by the presence of a magnetic field, but the increase in deformation with the different configurations of magnetic field was less pronounced, leading to nearly coincident curves.
 We present 3- and 2-dimensional illustrations of the NSs in Fig.
 4.
-We take examples from the case of $B_{0}=3\times 10^{18}$ G and $B_{surf}=10^{15}$ G.
+We take examples from the case of $\gamma=3$ with $\Omega=1000$ s-1.
 At lower mass, the star’s shape becomes very oblate.
-At $M=0.31\>M_{\odot}$, $R_{p}$ = 4.96 km, and $R_{e}$ = 12.51 km; at $M=0.51\>M_{\odot}$, $R_{p}$ = 7.75 km, and $R_{e}$ = 11.99 km; $M=1.18\>M_{\odot}$, $R_{p}$ = 10.67 km, and $R_{e}$ = 11.87 km; and $M=1.83\>M_{\odot}$, $R_{p}$ = 10.48 km, and $R_{e}$ = 10.63 km.
+At $M=0.9\>M_{\odot}$, $R_{POL}$ = 9.04 km, and $R_{EQ}$ = 11.12 km; at $M=1.4\>M_{\odot}$, $R_{POL}$ = 10.34 km, and $R_{EQ}$ = 11.32 km; at $M=1.7\>M_{\odot}$, $R_{POL}$ = 10.64 km, and $R_{EQ}$ = 11.11 km; and at $M=1.9\>M_{\odot}$, $R_{POL}$ = 10.31 km, and $R_{EQ}$ = 10.50 km.
 It is noteworthy to compare our results with those presented in Ref.
-[1].
+Pattersons2021 .
 The deformation of rotating NSs in our study is significantly more pronounced than in their findings, which are based on rotating NSs with anisotropic pressure of matter, but without considering the contribution of a magnetic field to the deformation.
 This fact strengthens our argument that the magnetic field strongly induces the deformation of rotating NSs.
+The oblateness of NSs can also be characterized by their eccentricity $e$.
+In the context of a two-dimensional body, when the value of $e$ approaches 0, the shape of the body tends to be circular.
+Conversely, as $e$ approaches 1, the shape becomes more elliptical.
+Fig.
+5 illustrates the relationship between mass and eccentricity.
+Both rotation and chaotic magnetic fields contribute to an increased deformation of NSs, as higher values of $\Omega$ and the appearance of $B$ drive the eccentricity $e$ closer to 1.
+Furthermore, it is noteworthy to analyze the trends of the eccentricity curves.
+At $\Omega=1000$ s-1, the overall trend is consistent with the results reported in Ref.
+Lopes2024 .
+As the rotational frequency increases, the trends become more comparable to those presented in Ref.
+Pattersons2021 .
+Fig.
+6 illustrates the relationship between the $M_{0}$ and the Kepler frequency of NSs.
+The Kepler frequency increases as the chaotic magnetic field appears, indicating that NSs with chaotic magnetic fields can rotate faster than those with weaker fields.
+Previously, the mass and radius calculations revealed that rotating NSs with chaotic magnetic fields exhibit lower maximum masses and smaller radii.
+This finding suggests that, for NSs rotating at their respective Kepler frequencies, our results are consistent with the conservation of angular momentum.
+Specifically, the presence of magnetic fields is associated with a reduction in both maximum mass and radius, but an increase in the maximum angular velocity (i.e., the Kepler frequency).
+These findings support the validity of the fact that the underlying physics captured by our results is accurate and physically consistent.
+Fig.
+7 depicts the relationship between the rotational mass and the moment of inertia of NSs and the corresponding eccentricity of the NSs.
+It is worth noting that the Eqs.
+(27) and (39) that help us to calculate the moment of inertia of NSs are only valid for spherical shape.
+So, to keep the validity of the spherical-body approximation, we set the angular velocity to $\Omega=50$ s-1.
+As shown in Fig.
+7(b), the resulting eccentricity values are found to be very close to zero, confirming that the stars remain nearly spherical under this condition.
+This justifies the continued use of the standard moment of inertia formula.
+Similar treatments for very slowly rotating NSs have also been employed in Refs.
+Rahmansyah2020 ; Rizaldy2024 , thereby providing precedent and support for our approach.
+Our motivation for retaining the investigation of the moment of inertia in this work is to obtain a qualitative insight into the effect of chaotic magnetic fields on the moment of inertia of NSs.
+Despite the simplification, this still provides a meaningful comparison with moment of inertia constraint, such as that reported in Ref.
+Landry2020 .
+Moreover, the study of the moment of inertia, even under the very slow rotation approximation, remains essential in neutron star physics, as it connects to various observable phenomena, such as pulsar glitches (e.g.
+see Ref.
+Haskell2015).
+Since magnetic fields can significantly modify the internal structure and moment of inertia, it is important to assess their qualitative impact, even if only within the slowly rotating approximation.
+We can see that at lower masses, the curves of moment of inertia tend to coincide, suggesting that the impact of chaotic magnetic field strength on the moment of inertia is negligible in this regime.
+However, a noticeable separation begins at $M=0.3$ $M_{\odot}$, where the appearance of chaotic magnetic fields lead to a reduction in the moment of inertia.
+Given that the moment of inertia depends on both the mass and the square of the radius, these results are consistent with the mass-radius relations.
+Remarkably, our findings for the moment of inertia are consistent with the constraint range provided by Landry et al.
+Landry2020 .
+It is worth noting that this constraint evaluates whether rotating NSs with chaotic magnetic fields align with observational evidence.
+The constraint is given as $I=1.51^{+0.20}_{-0.30}\times 10^{45}$ g cm2 at $M=1.4$ $M_{\odot}$, the standard canonical mass of NSs.
+As detailed in Ref.
+Landry2020 , this constraint is derived from measurements of pulsar masses, gravitational wave signals from events such as GW170817 and GW190425, and X-ray data from hotspot emissions on NS surfaces observed by NICER.
 4 Conclusion
@@
 Our findings indicate that the magnetic field can decrease the radius of NSs.
-NSs formed with stronger chaotic magnetic fields exhibit a lower maximum mass compared to those formed with weaker chaotic magnetic fields.
-Additionally, while the chaotic magnetic field can contribute to a reduction in radius, it can simultaneously enhance the compactness and deformation of rotating NSs, with significant deformation occurs at lower mass values.
-Acknowledgements
+NSs formed with chaotic magnetic fields exhibit a lower maximum mass compared to those without chaotic magnetic fields.
+Additionally, while the chaotic magnetic field can contribute to a reduction in radius, it can simultaneously enhance the compactness and deformation of rotating NSs.
+For deformed rotating NSs, significant deformations occur at lower mass values.
+On the other side, our results indicate that the presence of chaotic magnetic field enhances the Kepler frequency of rotating NSs, whereas it simultaneously tend to decrease their moment of inertia.
+Additionally, the moment of inertia of rotating NSs in our study is consistent with the constraint range obtained from pulsar mass measurements, gravitational wave event data (GW170817 and GW190425), and X-ray observations of emissions from hotspots on NS surfaces measured by NICER.
+Acknowledgements.
+We sincerely thank Tsuyoshi Miyatsu and Ryan Rizaldy for precious discussions.
+Special thanks are given to Anna Campoy Ordaz for making her code publicly available, which allowed MLP to modify it for use in this work.
 MLP acknowledges the Indonesia Endowment Fund for Education Agency (LPDP) for financial support.
-Special thanks are extended to Anna Campoy Ordaz for making her code publicly available, which allowed MLP to modify it for use in this work.
-FPZ and GH would like to thank Kemendikbudristek Republic of Indonesia (through LPPM ITB, Indonesia) for partially financial support.
+FPZ and GH would like to thank Kemendiktisaintek (The Ministry of Higher Education, Science, and Technology) of the Republic of Indonesia (through LPPM ITB, Indonesia) for partially financial support.
+HLP and MLP would like to thank the members of the Theoretical Physics Groups at Institut Teknologi Bandung for their hospitality.
+MFARS is supported by the Second Century Fund (C2F), Chulalongkorn University, Thailand.